OpenStudy (anonymous):

The formula V = IR shows the relationship between voltage, current, and resistance. Solve this formula for R. R = VI R = V divided by I R = V + I R = V – I

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yes it B Correct wee thats what i git to

OpenStudy (anonymous):

Marcie wants to enclose her yard with a fence. Her yard is in the shape of a rectangle attached to a triangle. The formula for the area of the enclosed space is A = lw + 0.5bh. Solve for b. b = A divided by the quantity l times w plus 0.5 times h b = The quantity A minus l times w all divided by 0.5 times h b = A – lw – 0.5h b = A + lw + 0.5h

OpenStudy (johnweldon1993):

\[\large A = LW + 0.5bh\] subtract LW from both sides of the equation \[\large A - LW = 0.5bh\] Divide both sides by 0.5H \[\large \frac{A - LW}{0.5H} = b\]